OpenStudy (anonymous):

-3x+2y=2 all i need is one more point .... help???

OpenStudy (primeralph):

@robz8 Am I the only one getting confused with these questions?

OpenStudy (anonymous):

like i said before, just pick a number for x and see what y equals order pairs are in the form (x,y)

OpenStudy (primeralph):

I don't know if it should be harder than it see,s/

OpenStudy (primeralph):

*seems

OpenStudy (anonymous):

i did but i keep getting it rong and imma bout to throw some pellet

OpenStudy (primeralph):

You meant "sh*t" right?

OpenStudy (anonymous):

not sure how you could get it wrong?

OpenStudy (anonymous):

yeah thats wat i typed and it changed it

OpenStudy (anonymous):

and idk either

OpenStudy (anonymous):

pick a random number and tell it to me

OpenStudy (anonymous):

69 lmafo jkjk umm 5

OpenStudy (anonymous):

-3*(5) + 2y = 2 -15 + 2y = 2 2y = 17 y = 8.5 (5, 8.5) is your point

OpenStudy (anonymous):

i think x is -2 and y is also -2

OpenStudy (anonymous):

but i cant plot a decimal on a graph

OpenStudy (anonymous):

8.5 = 17/2 or you could just pick a different number for x...

OpenStudy (anonymous):

8??

OpenStudy (anonymous):

okay, plug 8 where x is in the equation and solve for y (8, y) will be your point

OpenStudy (primeralph):